can you help me solve this problem:
y=2x+7
2x-y=-15
i know u have to do the substitution method but i tried so many times. thanks 4 ur help

Answer by jim_thompson5910(35256)

y = 2x+7

so y and 2x+7 are the same thing

this means you can replace all/any y values in the second equation with 2x+7

to go from

2x-y=-15

to

2x-(2x+7)=-15

Then distribute to get

2x - 2x - 7 = -15

and combine like terms to get

0x - 7 = -15

Then simplify to get

-7 = -15

This is NEVER true regardless of what the value of x or y is.

So there are NO solutions to this system because the system is inconsistent